OpenStudy (anonymous):

Compute the distance Erica will run if she continues using the formula, y = 3(1.1)x, for 52 weeks (one year). Does this distance seem reasonable? A. yes B. no i put A (: correct?

OpenStudy (anonymous):

Well, if x is measured in weeks, that is 3*1.1^52 = 3 * 142 = 426 miles. I guess it depends on your definition of reasonability, but I would say...No!! :)

OpenStudy (anonymous):

Alrighty Thanks!

OpenStudy (dumbcow):

i dont know for a whole year i think its reasonable it comes out to about 1.2 miles per day some people do that...not me though